OpenStudy (amy0799):

consider the following. f(x) = (x)/(x^2-2) Find the x-values at which f is not continuous, smaller value = larger value = Is the discontinuity at the smaller value removable? Is the discontinuity at the larger value removable?

OpenStudy (anonymous):

look for a value of x that makes the denominator zero, it won't be continuous there

OpenStudy (michele_laino):

hint: for continuity, we can not divide by zero

OpenStudy (amy0799):

i'm confused on what to do...

OpenStudy (michele_laino):

we have tosolve this equation, first: \[\Large {x^2} - 2 = 0\]

OpenStudy (amy0799):

x=sqrt(2)

OpenStudy (michele_laino):

more precisely, we have two solutions: \[\Large x = \pm \sqrt 2 \]

OpenStudy (michele_laino):

am I right?

OpenStudy (amy0799):

yea. so the smaller value is \[-\sqrt{2}\] and the larger value is \[\sqrt{2} \]?

OpenStudy (michele_laino):

Correct! we have to exclude both values, since we can not divide by zero, so our function is not continuous at these points: \[\Large \begin{gathered} x = \sqrt 2 \hfill \\ x = - \sqrt 2 \hfill \\ \end{gathered} \]

OpenStudy (amy0799):

right. so how do you find out if it's removable?

OpenStudy (michele_laino):

I think that both discontinuities are not removable

OpenStudy (amy0799):

can you explain why?

OpenStudy (michele_laino):

yes! I explain with an example

OpenStudy (michele_laino):

for example, the subsequent function: \[\Large f\left( x \right) = \frac{{{x^2} - 1}}{{x - 1}}\] is not continuous at x=1, right?

OpenStudy (amy0799):

yes

OpenStudy (michele_laino):

nevertheless x=1 is a removable discontinuity, since we can write this: \[\Large f\left( x \right) = \frac{{{x^2} - 1}}{{x - 1}} = \frac{{\left( {x - 1} \right)\left( {x + 1} \right)}}{{x - 1}} = x + 1\]

OpenStudy (amy0799):

ooh ok. can you help me with another problem?

OpenStudy (michele_laino):

yes!

OpenStudy (amy0799):

consider the following. f(x) = {-9x

OpenStudy (michele_laino):

ok!

OpenStudy (amy0799):

f(x) = -9 if x less than or equal to 5 x^2-2x+2 if x greater than 5

OpenStudy (amy0799):

find the x-value at which f is not continuous and is the discontinuity removable

OpenStudy (michele_laino):

at x= 5, we have f(x)= 9 using your first part for f(x) and we have: \[{x^2} - 2x + 2 = 25 - 2 \cdot 5 + 2 = 17 \ne 9\] so the discontinuity is at x=5

OpenStudy (amy0799):

is it not removable?

OpenStudy (michele_laino):

I think that it is not removable

OpenStudy (michele_laino):

because it is a jump-type discontinuity

OpenStudy (amy0799):

ok thanks!. I have a few more I need help me, is that ok?

OpenStudy (michele_laino):

ok!

OpenStudy (amy0799):

g(x) = 5sin(x)/x if x less than 0 a-4x if x greater than or less than 0 Find the constant a such that the function is continuous on the entire real line.

OpenStudy (michele_laino):

we can write this: \[\Large \mathop {\lim }\limits_{x \to 0} \frac{{5\sin x}}{x} = 5\]

OpenStudy (michele_laino):

whereas at x=0, we have: \[\Large a - 4x = a - 4 \cdot 0 = a\]

OpenStudy (michele_laino):

now, for continuity, we have to request taht both values are equal each to other. So what do you get?

OpenStudy (amy0799):

Would the constant a be 5?

OpenStudy (michele_laino):

that*

OpenStudy (michele_laino):

correct! a=5

OpenStudy (amy0799):

g(x) = (x^2-a^2)/(x-a) if x is not equal to a 11 if x=a Find the constant a such that the function is continuous on the entire real line.

OpenStudy (michele_laino):

we can write this: \[\Large \frac{{{x^2} - {a^2}}}{{x - a}} = \frac{{\left( {x - a} \right)\left( {x + a} \right)}}{{x - a}} = x + a\]

OpenStudy (michele_laino):

and its value at x=a, is: \[\Large 2a\] Now, for continuity we have to request, as before, that those two values are equal each to other, so we have to solve this equation: \[\Large 2a = 11\]

OpenStudy (amy0799):

I don't understand how you got 2a

OpenStudy (michele_laino):

after a simplification, I got this: \[\Large g\left( x \right) = x + a\] so, \[\Large g\left( a \right) = a + a = 2a\]

OpenStudy (amy0799):

oh ok. so the constant a is 11/2?

OpenStudy (michele_laino):

yes! correct!

OpenStudy (amy0799):

Describe the interval(s) on which the function below is continuous. \[f(x)=x \sqrt{x+1}\]

OpenStudy (michele_laino):

your function is continuous inside all its domain, and its domain is the subset of all real numbers, such that the radicand is positive or equal to zero, namely: \[\Large x + 1 \geqslant 0\]

OpenStudy (amy0799):

[0,∞) is that how you would write it?

OpenStudy (michele_laino):

we have: \[\Large x + 1 \geqslant 0 \Rightarrow x \geqslant - 1\] so, the requested interval is: \[\Large [ - 1,\infty )\] am I right?

OpenStudy (amy0799):

yes. Thank you so much for the help and sorry if I took up your time.

OpenStudy (michele_laino):

no worries! :) :)